OpenStudy (anonymous):

solve the equation by using the square root property simplify all radicals 2x^2-27=351

OpenStudy (cwrw238):

add 27 to both sides 2 x^2 = 351 + 27 = 378 x^2 = 378/2 = 189 x = sqrt 189

OpenStudy (cwrw238):

use your calculator to find this

OpenStudy (anonymous):

@cwrw238 got it

OpenStudy (anonymous):

3x^2=5

OpenStudy (anonymous):

@cwrw238 for that last answer i did find the radical for it